d2^x/dt^2 - 6 dx/dt +9x =0; when t =0, x =2 and dx/dt = 0
1
Expert's answer
2020-10-05T19:05:29-0400

"\frac{d^2x}{dt^2}-6\frac{dx}{dt}+9x=0", "x(0)=2" and "\frac{dx(0)}{dt}=0".


Let us solve the characteristic equation:


"k^2-6k+9=0"

"(k-3)^2=0"

"k_1=k_2=3"


Therefore, the general solution is of the following form:


"x(t)=e^{3t}(C_1+C_2t)"


Then "\frac{dx}{dt}=3e^{3t}(C_1+C_2t)+C_2e^{3t}=e^{3t}(3C_1+3C_2t+C_2)"


Taking into account that "x(0)=2" and "\frac{dx(0)}{dt}=0" we conclude that "C_1=2" and "3C_1+C_2=0". Thus "C_2=-3C_1=-6".


Consequently, the particular solution is the following:


"x(t)=e^{3t}(2-6t)"